# Initialization Is Critical to Whether Transformer fits Composite Function by Inference or Memorizing

### v1

@@ 5 Mechanisms of Models in Two Phases
In this section, we analyze mechanisms underlying different learning solutions including: i) information flow via attention matrix; ii) information representation in vector space; iii) model complexity form different perspectives. A key message is as follows. A model with small initialization needs to gradually increase its capacity by enlarging the parameter scale, therefore, to use the least optimization cost to fit the training data, the model only needs to learn four elementary anchor functions to obtain inferential mappings. As the initialization scale gets larger, it tends to learn ten symmetric mappings, and memorize the mapping between anchor-key item pairs and their outputs.
@@ 5.1 Information Transmission and Fusion Mechanisms
In this section, we study the information transmission and fusion mechanisms occurring in the attention matrix through information flow analysis. As shown in Fig. 3 (a,c), we present the information flow of two-layer networks corresponding to two solutions. We use the same input sequence $X=(35,63,99,4,3,38,42,36,62)$ to test the output of the key item $x_{3}=99$ under the influence of the composite anchor pair $(x_{4},x_{5})=(4,3)$. The thickness of the line connecting the $j$-th token in Layer $l$ and the $k$-th token in Layer $l+1$ represents the value of the attention matrix $\mathrm{Attn}^{(l)}$ at position ($k$, $j$). For ease of observation, we manually annotate the information flow that significantly contributes to the output of the last token using different colors.
@@
Figure 3: (a, c) Information flow in the two-layer networks of symmetric and inferential solutions. The input sequence shown in the figure represents the test sample, with key items and anchor positions annotated. For each layer’s attention matrix, we illustrate the mechanisms of information transmission and fusion through the information flow. The thickness of the line represents the value of the attention matrix $\mathrm{Attn}^{(l)}$. We use different colors to mark the key item and the two single anchors, and highlight the attention connections that significantly contribute to the final output. The final output sequence represents the model’s output. (a) Symmetric solution. (c) Inferential solution. (b) T-SNE visualization of vectors $X^{\mathrm{ao}(1)}$ of 10,000 input sequences with different anchor-key item pairs.
@@
For symmetric solutions, we find that the model’s attention mechanism plays a key role in enabling the learning of consistent mappings for symmetric anchor pairs. In the first layer of the network, the attention mechanism combines the information from the two anchors and places it in the last token of the sequence. This allows the model to learn identical embeddings for symmetric anchor-key pairs, such as (3, 4) and (4, 3), as illustrated in Fig. 3(a). In the second layer, the information of the key item is broadcast to the last token and combined with the anchor information obtained from the first layer by the residual connection. This enables the model to produce the final output based on the combined information from the anchors and the key item.
@@
To verify the symmetric nature of the learned representations after the first-layer attention, we visualize the vectors $X^{\mathrm{ao}(1)}$ of 10,000 input sequences with different anchor-key item pairs using t-SNE. Fig. 3(b) shows that symmetric anchor-key pairs (e.g., (3, 4) and (4, 3)) are clustered together in the $X^{\mathrm{ao}(1)}$ vector space, confirming that the model learns to map them to similar representations.
@@
For inferential solutions, the information transmission mechanism is different. As shown in Fig. 3(c), in the first layer, the key item information is moved to the positions of the two anchors, and each anchor is combined with the key item information separately. This allows the model to learn individual mappings for each single anchor. In the second layer, the anchor tokens (now containing the combined information from the anchors and the key item) are broadcast to the last token and combined to produce the final output. This composition of single anchor mappings enables the model to learn inferential solutions that generalize to unseen composite anchors.
@@ 5.2 Divergence in Fused Vector Representations across Two Phases
To further investigate the mechanistic differences between symmetric and inferential solutions, we examine the divergence in vector representations after the fusion of anchor and key item information. Specifically, for different anchor-key item combinations, we explore the cosine similarity between the output vectors $X^{\mathrm{ao}(2)}_{[9,:]}$, i.e., the last token of the output vectors of the second attention layer, under the two mechanisms. Let $\bm{v}_{(x;a_{1},a_{2})}$ denote the output vector $X^{\mathrm{ao}(2)}_{[9,:]}$ for a sequence with key item $x$ and anchors $(a_{1},a_{2})$ [^1] For simplicity, we assume that the position of the anchor-key item combination and other items do not significantly affect the output vector..
@@
As shown in Fig. 4 (a, b), the heatmaps display the cosine similarity between output vectors for different key item and anchor combinations. Both the $x$-axis and $y$-axis represent the values of key items, ranging from 30 to 40. The lower triangle of the heatmap corresponds to the cosine similarity matrix between $\bm{v}_{(x;3,3)}$ and $\bm{v}_{(y;2,2)}$, where $x$ and $y$ are the key items at the corresponding positions on the $x$-axis and $y$-axis, respectively. Similarly, the upper triangle matrix represents the similarity matrix between $\bm{v}_{(x;1,2)}$ and $\bm{v}_{(y;1,3)}$. The red boxes highlight the positions where the inferential targets are the same, for example, the red highlight in the first column in Fig. 4(a) indicates $f(36;3,3)=f(30;2,2)$.
@@
For inferential solutions, and for any key items $x_{1}$, $x_{2}$, anchors $(a_{1,1},a_{1,2})$, $(a_{2,1},a_{2,2})$, if $f(x_{1};a_{1,1},a_{1,2})=f(x_{2};a_{2,1},a_{2,2})$, then the vectors $\bm{v}_{(x_{1};a_{1,1},a_{1,2})}$ and $\bm{v}_{(x_{2};a_{2,1},a_{2,2})}$ are nearly parallel. In contrast, for symmetric solutions, the pairwise similarity between vectors $\bm{v}_{(x;a_{1},a_{2})}$ is always low even for those with the same output. This indicates that the last FNN layer of the symmetric solution memorize all possible projections from $X^{\mathrm{ao}(2)}_{[9,:]}$ to the output of the decoder layer $X^{\mathrm{do}(2)}_{[9,:]}$, while the last FNN for the inferential solution only needs to learn much fewer projections.
@@
Figure 4: Cosine similarity heatmaps. Each axis represents a selected anchor, with the value on the coordinate axis representing the value of the key item. Color indicates the cosine similarity between specific vectors, which are defined within the respective subplot. The red boxes highlight the positions where the target outputs obtained by the anchors on the $x$-axis and $y$-axis are the same for the corresponding key items. (a, b) Color indicates the cosine similarity between the last token of $X^{\mathrm{ao}(2)}$ of different anchor-key item pairs. (a) Inferential solution. (b) Symmetric solution. (c) Color indicates the cosine similarity between $V^{(2)}$ rows corresponding to the first anchor’s position across different anchor-key pairs for inferential solutions.
@@
The core of learning inferential solutions lies in the model’s ability to infer the operation of each single anchor. To further verify the alignment between the single anchor operations in the inferential mechanism and the defined single-anchor function $g$, we investigate the vector representations after the fusion of the first anchor and the key item. From the information flow diagram (Fig. 3(c)), the information from the two anchors and the key item is fused through the attention mechanism in the second layer. Therefore, we focus on the vector representation of the second layer Value matrix $V^{(2)}$ at the position of the first single anchor. Similar to the approach in Fig. 4(a, b), we examine the structural characteristics of the output $\bm{u}_{(x;a_{1})}$ at the position of the first anchor of $V^{(2)}$ for the sequence with key item $x$ and the first anchor $a_{1}$. The red boxes highlight the positions where the targets obtained by the single anchors $a_{1}$ and $a_{2}$ on the $x$-axis and $y$-axis are the same for the corresponding key items $x_{1}$ and $x_{2}$, i.e., $g(x_{1};a_{1})=g(x_{2};a_{2})$. As shown in Fig. 4(c), for single anchor mappings, the information fusion process of single anchors and key items during the inferential solution’s information fusion and the explicitly defined single anchor operations align well. This provides stronger evidence for the model’s ability to learn inferential solutions.
@@ 5.3 Model Complexity: A Key Factor in Phase Transitions
Large initialization scale endows the model with large complexity with training data with changing parameters slightly, such as the case of fully-connected neural networks (FNN) in linear regime (e.g., Neural tangent kernel) (Luo et al., 2021; Jacot et al., 2018). However, the model with small initialization with inadequate complexity often gradually increases its complexity during the training. For example, neurons in the same layer of FNNs with small initialization would first condense in very few directions (Zhou et al., 2021), followed by condensing in more and more directions (Luo et al., 2021; Zhang et al., 2021, 2022), which can regularize FNNs to fit data with as few condensation directions as possible. Therefore, we investigate the degree of parameter condensation in models with inferential or symmetric solutions.
Figure 5: (a, b) Cosine similarity of neurons in the $W^{Q(1)}$ matrix. (a) Inferential solution with small initialization. (b) Symmetric solution with large initialization. (c, d) Visualization of the embedding space using PCA for different initialization scales. (c) Inferential solution with small initialization. The embedded tokens seem to form arithmetic sequences with common differences of 3 (red arrow) and 4 (blue arrow) along the two directions. (d) Symmetric solution with large initialization.
@@
Structural Features of the Word Embedding Matrix. To gain further insights into the learning mechanisms of inferential solutions, we analyze the structure of the model’s embedding matrix. We visualize the embedding space using Principal Component Analysis (PCA). As shown in Fig. 5(c,d), for small initialization scales (Fig. 5(c)), we observe a clear ordinal structure in the embedding space, with the embeddings of the input tokens arranged according to their numerical value. This also suggests a low-complexity tendency in the word embedding matrix, requiring the model to capture the relative ordinal relationships between different tokens. However, this ordinal structure is not present for large initialization scales (Fig. 5(d)).
It is worth noting that the relative ordinal relationship in the inferential solution is not a simple numerical magnitude relationship of the corresponding tokens. We observe that this ordinal relationship may originate from the definition of the four single anchors, where the differences between the operations of any two single anchors can be obtained by the addition of the basic elements 3 and 4. This arrangement leads to the numbers being ordered with intervals of 3 (red arrow) and 4 (blue arrow) from two directions in the embedding space. To further verify the low complexity bias of word embedding matrices, we visualize the eigenvalues of the covariance matrix of the embedding vectors, the results are shown in Appendix B.
@@ 6 Prediction of mechanisms underlying different solutions
As the data complexity increases, neural networks usually need more training steps to fit the data (Arpit et al., 2017; Xu et al., 2019, 2020; Rahaman et al., 2019). We define the number of composite anchor pair groups [^2] Two anchor pairs with symmetrical relationships are considered to be in the same group. that do not satisfy the inferential mapping as the inferential complexity.
It is anticipated that the model of small initialization, i.e., fitting data with as low complexity as possible, should use more training steps to fit data with larger inferential complexity, while the model with large initialization, i.e., easily memorizing ten mappings of the symmetric solution, should use similar training steps to fit data with different inferential complexity.
Figure 6: Training speed with different initialization scales when facing different data complexity. Abscissa: the model depth. Ordinate: the number of epochs required for the model’s training loss to reach $5\times 10^{-2}$. Color: the number of noise groups. We conduct 9 random trials for each setting of different initializations, depths, and data inferential complexities, and take the average of the required number of epochs. (a) Small initialization. (b) Large initialization.
To measure the training steps required to fit data, we use the number of epochs required for the model’s training loss to reach $5\times 10^{-2}$. We conduct 9 random trials for each setting. As shown in Fig. 6, models that tend to learn data with inferential solutions indeed use more training steps for data with larger inferential complexity (Fig. 6(a)), while models that tend to learn data with symmetric solutions use roughly the same training steps for data with different inferential complexity (Fig. 6(b)). This prediction convincingly backs up our analysis of why different initialization scales can lead to different solutions.
@@ 7 Discussion
In this work, we investigate the influence of parameter initialization scale on the solutions learned by transformers for compositional tasks. We discover a phase diagram of model solutions, where small initialization scales lead to inferential solutions that capture the underlying compositional primitives, while larger initialization scales result in symmetric solutions that memorize mappings. To explain these findings, we analyze the information flow and vector representations within the model, revealing distinct mechanisms for inferential and symmetric solutions. Inferential solutions exhibit low complexity and learn individual mappings for single anchors, while symmetric solutions directly combine anchor information without learning the compositional structure. Building upon these insights, we can predict and experimentally confirm that models with small initialization scales require more training steps to fit data with larger inferential complexity, while models with large initialization scales are less affected by inferential complexity.
Our work opens up promising avenues for future research, such as leveraging Mixture of Experts (MoE) to design networks with different initialization scales for different expert models, and exploring improved training strategies that adapt the model’s focus on enhancing memorization or inference capabilities at different stages of training. It is important to acknowledge the limitations of our current study. The experiments and analyses are based on synthetic data, which allows us to cleanly isolate and characterize the observed phenomena. However, the applicability of our findings to more realistic and complex datasets remains to be validated. In future work, we plan to extend our investigation to real-world datasets and tasks, aiming to bridge the gap between our theoretical understanding and the practical application of these models in solving compositional problems.

### v2

@@ 5 Mechanisms of Models in Two Phases
In this section, we analyze mechanisms underlying different learning solutions including: i) information flow via attention matrix; ii) information representation in vector space; iii) model complexity from different perspectives. The key message is as follows. A model with small initialization needs to gradually increase its capacity by enlarging the parameter scale, therefore, to use the least optimization cost to fit the training data, the model only needs to learn four single anchor functions to obtain inferential mappings. As the initialization scale increases, the model tends to learn ten mappings, treating symmetric anchor pairs as equivalent, and memorizes the mapping between each anchor-key item pair and its corresponding output. When the initialization becomes even larger, the model forgoes learning general patterns and instead merely memorizes the output associated with each individual data point. This causes the model to lose generalization ability, even on the anchors it has encountered during training.
@@ 5.1 Information Transmission and Fusion Mechanisms
In this section, we study the information transmission and fusion mechanisms occurring in the attention matrix through information flow analysis. As shown in Fig. 3 (a,c), we present the information flow of two-layer networks corresponding to the two solutions. We use the same input sequence to test the output with the key item $99$ and the unseen anchor pair $(4,3)$. The thickness of the line connecting the $j$-th token in Layer $l$ and the $k$-th token in Layer $l+1$ represents the value of the attention matrix $\mathrm{Attn}^{(l)}$ at position ($k$, $j$). For ease of observation, we manually annotate the information flow that significantly contributes to the output of the last token using different colors.
@@
Figure 3: (a, c) Information flow in the two-layer networks of symmetric and inferential solutions. The input sequence shown in the figure represents the test sample, with key items and anchor positions annotated. For each layer’s attention matrix, we illustrate the mechanisms of information transmission and fusion through the information flow. The thickness of the line represents the corresponding value in the attention matrix $\mathrm{Attn}^{(l)}$. We use different colors to mark the key item and the two single anchors, and highlight the attention connections that significantly contribute to the final output. The final output sequence represents the model’s output. (a) Symmetric solution. (c) Inferential solution. (b) T-SNE visualization of vectors $X^{\mathrm{ao}(1)}$ of 10,000 input sequences with different anchor-key item pairs. Symmetric anchor pairs have similar colors in different shades.
@@
For symmetric solutions, we find that the model’s attention mechanism plays a key role in enabling the learning of consistent mappings for symmetric anchor pairs. In the first layer of the network, the attention mechanism combines the information of the two anchors and moves it to the last token of the sequence. This allows the model to learn identical embeddings for symmetric anchor pairs, such as (3, 4) and (4, 3), as illustrated in Fig. 3(a). In the second layer, the information of the key item is broadcast to the last token and combined with the anchor information obtained from the first layer by the residual connection. This enables the model to produce the final output based on the combined information from the anchors and the key item.
@@
To verify the symmetric nature of the learned representations after the first-layer attention, we visualize the vectors $X^{\mathrm{ao}(1)}$ of 10,000 input sequences with different anchor-key item pairs using t-SNE. Fig. 3(b) shows that symmetric anchor pairs are clustered together in the $X^{\mathrm{ao}(1)}$ vector space, confirming that the model learns to map them to similar representations.
@@
For inferential solutions, the information transmission mechanism is different. As shown in Fig. 3(c), in the first layer, the key item information is moved to the positions of the two single anchors, and each anchor is combined with the key item information separately. This allows the model to learn individual mappings for each single anchor. In the second layer, the anchor tokens (now containing the combined information from the anchors and the key item) are moved to the last token and combined to produce the final output. This combination mechanism enables the model to learn inferential solutions on the composite anchors.
@@ 5.2 Divergence in Fused Vector Representations across Two Phases
To further investigate the mechanistic differences between symmetric and inferential solutions, we examine the divergence in vector representations after the fusion of anchor pair and key item information in both types of solutions. Specifically, we examine the cosine similarity between the output vectors of the second attention layer’s last token (i.e., the last token of $X^{\mathrm{ao}(2)}$) for different anchor-key item combinations. We denote these output vectors as $\bm{v}(x;a_{1},a_{2})$, where $x$ is the key item and $(a_{1},a_{2})$ is the anchor pair of the input sequence [^1] For simplicity, we assume that the position of the anchor-key item combination and other items do not significantly affect the output vector..
@@
As illustrated in Fig. 4(a, b), the heatmaps display the cosine similarity between output vectors for various anchor-key item pairs. Both the abscissa and ordinate represent key item values ranging from 30 to 40. The lower and upper triangles of the heatmap correspond to the cosine similarity matrices between $\bm{v}(x_{1};3,3)$ and $\bm{v}(x_{2};2,2)$, and between $\bm{v}(x_{1};1,2)$ and $\bm{v}(x_{2};1,3)$, respectively, where $x_{1}$ and $x_{2}$ are the corresponding key item values on the axes. The red boxes highlight positions where the inferential targets are the same, for example, the red highlight in the first column in Fig. 4(a) indicates $f(36;3,3)=f(30;2,2)$, where $f(\cdot;\cdot,\cdot)$ is defined in Equation 1.
@@
For inferential solutions, if two anchor-key pairs $(x_{1};a_{1},a_{2})$ and $(x_{2};b_{1},b_{2})$ have the same output, i.e., $f(x_{1};a_{1},a_{2})=f(x_{2};b_{1},b_{2})$, then their fused vectors $\bm{v}(x_{1};a_{1},a_{2})$ and $\bm{v}(x_{2};b_{1},b_{2})$ are nearly parallel. In contrast, for symmetric solutions, the pairwise similarity between vectors $\bm{v}(x;a_{1},a_{2})$ is always low even for those with the same output. This suggests that the last FNN layer of the symmetric solution memorizes all possible projections from $X^{\mathrm{ao}(2)}$ to the decoder layer output. Conversely, the last FNN of the inferential solution only needs to learn fewer projections, as the same output is represented by similar vectors in the $X^{\mathrm{ao}(2)}$ space.
@@
Figure 4: Cosine similarity heatmaps for vector representations in different solutions. Each axis represents a selected anchor pair (labeled on the axis), with the value on the coordinate axis representing the value of the key item.The color indicates the cosine similarity between specific vectors defined in each subplot. Red boxes highlight positions where the target outputs obtained by the anchors on the abscissa and ordinate are the same for the corresponding key items. (a, b) Cosine similarity between the output vectors of the second attention layer’s last token (the last token of $X^{\mathrm{ao}(2)}$) for different anchor-key item pairs in (a) inferential and (b) symmetric solutions. (c) Cosine similarity between the rows of the second layer Value matrix ($V^{(2)}$) corresponding to the first anchor’s position across different anchor-key item pairs for inferential solutions.
@@
Learning inferential solutions relies on the model’s ability to infer the operation of each individual anchor. To verify the alignment between the single anchor operations in the inferential mechanism and the defined single-anchor function $g(\cdot;\cdot)$ in Equation (2), we investigate the vector representations after the fusion of the first single anchor and the key item. We focus on the vector representation of the second layer Value matrix $V^{(2)}$ at the position of the first single anchor, denoted as $\bm{u}(x;a_{1})$, with key item $x$ and first anchor $a_{1}$. Similar to Fig. 4(a, b), the red boxes highlight the positions where the targets obtained by the single anchors $a_{1}$ and $a_{2}$ on the abscissa and ordinate are the same for the corresponding key items $x_{1}$ and $x_{2}$, i.e., $g(x_{1};a_{1})=g(x_{2};a_{2})$. As shown in Fig. 4(c), the information fusion of single anchors and key items and the explicitly defined single anchor operations align well. This provides stronger evidence for the model’s ability to learn the mapping of each single anchor.
@@ 5.3 Model Complexity: A Key Factor in Phase Transitions
Large initialization scales endow models with high complexity, allowing them to fit training data with minor parameter changes, as seen in FNNs in the linear regime (e.g., Neural tangent kernel) (Jacot et al., 2018; Luo et al., 2021). Conversely, models with small initialization scales start with low complexity and gradually increase it during training. In FNNs with small initialization, neurons in the same layer first condense in a few directions (Zhou et al., 2021), i.e., their input weights align and concentrate in a few directions, followed by condensing in more directions (Luo et al., 2021; Zhang et al., 2022, 2021), regularizing FNNs to fit data with minimal condensation directions. This phenomenon of parameter condensation is closely related to the model’s complexity and its ability to learn the underlying structure of the data. To better understand the mechanisms behind inferential and symmetric solutions, we investigate the degree of parameter condensation.
Figure 5: (a, b) Cosine similarity of neurons in the $W^{Q(1)}$ matrix. (a) Inferential solution with small initialization. (b) Symmetric solution with large initialization. (c, d) Visualization of the embedding space using T-SNE for different initialization scales. (c) Inferential solution with small initialization. The embedded tokens seem to form arithmetic sequences with common differences of 3 (red arrow) and 4 (blue arrow) along the two directions. (d) Symmetric solution with large initialization. Please refer to Appendix D for more detailed experimental results under different model depths and initialization rates $\gamma$.
@@
Structural Features of the Word Embedding Matrix. To gain further insights into the learning mechanisms of inferential solutions, we analyze the structure of the model’s embedding matrix. We visualize the embedding space using T-SNE. As shown in Fig. 5(c,d), for small initialization scales (Fig. 5(c)), we observe a clear ordinal structure in the embedding space, with the embeddings of the input tokens arranged according to their numerical value. This also suggests a low-complexity tendency in the word embedding matrix, requiring the model to capture the relative ordinal relationships between different tokens. However, this ordinal structure is not present for large initialization scales (Fig. 5(d)).
It is worth noting that the relative ordinal relationship in the inferential solution is not a simple numerical magnitude relationship of the corresponding tokens. We observe that this ordinal relationship may originate from the definition of the four single anchors, where the differences between the operations of any two single anchors can be obtained by the addition of the basic elements 3 and 4. This arrangement is consistent with the numbers being ordered with intervals of 3 (red arrow) and 4 (blue arrow) from two directions in the embedding space. To further verify the low complexity bias of the word embedding matrix, we visualize the eigenvalues of the covariance matrix of the embedding vectors and its evolution process, the results are shown in Appendix D.3.
@@ 6 Prediction of mechanisms underlying different solutions
As the data complexity increases, neural networks usually need more training steps to fit the data (Arpit et al., 2017; Rahaman et al., 2019; Xu et al., 2020, 2019). Based on our analysis of the mechanisms underlying different solutions, we can predict that models with different initialization scales will exhibit varying degrees of difficulty in learning data with different complexities. We define the inferential complexity as the number of anchor pair groups [^2] Two anchor pairs with symmetrical relationships are considered to be in the same group. that do not satisfy the inferential mapping.
It is anticipated that the model of small initialization, i.e., fitting data with as low complexity as possible, should use more training steps to fit data with larger inferential complexity, while the model with large initialization, i.e., easily memorizing the mappings of the symmetric solution, should use similar training steps to fit data with different inferential complexity.
Figure 6: Training speed with different initialization scales when facing different data complexity. Abscissa: the model depth. Ordinate: the number of epochs required for the model’s training loss to reach $5\times 10^{-2}$. Color: the inferential complexity. We conduct 9 random trials for each setting of different initializations, depths, and data inferential complexities, and take the average of the required number of epochs. (a) Small initialization. (b) Large initialization.
To measure the training steps required to fit data, we use the number of epochs required for the model’s training loss to reach $5\times 10^{-2}$. We conduct 9 random trials for each setting. As shown in Fig. 6, models that tend to learn data with inferential solutions indeed use more training steps for data with larger inferential complexity (Fig. 6(a)), while models that tend to learn data with symmetric solutions use roughly the same training steps for data with different inferential complexity (Fig. 6(b)). This prediction convincingly backs up our analysis of why different initialization scales can lead to different solutions. We further demonstrate the learning speed of each anchor pair in Appendix F.
@@ 7 Discussion
Conclusion. In this work, we investigate the influence of parameter initialization scale on the solutions learned by transformers for compositional tasks. We discover a phase diagram of model solutions, where small initialization scales lead to inferential solutions that capture the underlying compositional primitives, while larger initialization scales result in symmetric solutions that memorize mappings. To explain these findings, we analyze the information flow and vector representations within the model, revealing distinct mechanisms for inferential and symmetric solutions. Inferential solutions exhibit low complexity and learn individual mappings for single anchors, while symmetric solutions directly combine anchor information without learning the compositional structure. Building upon these insights, we can predict and experimentally confirm the differing effects of inferential complexity on the training speed of models with different initialization scales.
Limitation and Future Work. The key limitation of our work is that the experiments and analyses are based on synthetic data, which may not fully capture the complexities of real-world datasets and tasks. Although some of our conclusions have also been validated in models like GPT-2, further verification on a wider range of data and models is necessary. Going forward, we plan to extend our investigation to real-world datasets and tasks, to bridge the gap between our theoretical understanding and practical application. This could involve leveraging Mixture of Experts (MoE) to design networks with different initialization scales for different expert models, and exploring improved training strategies that adapt the model’s focus on enhancing memorization or inference capabilities at different stages of training with different initialization scales.